### v1

Abstract
Cosmological gravitational particle production (GPP) is a generic mechanism by which particles are produced during the inflationary epoch. In this work we consider the GPP of massive scalars in an effort to fully understand the spectrum of the produced particles. We consider scalars which are conformally or minimally coupled to gravity, as well as models of both early and late reheating. We numerically calculate the particle production in each scenario and compare the results to analytic approximations from boundary matching, Stokes phenomenon, steepest descent, and inflaton scattering methods. For each method, we describe the regime of validity and show that there is good agreement between the analytic and numerical results.
1 Introduction
It is well known that particle production can occur in the early universe as a result of the rapid expansion of spacetime during inflation. This effect is known as cosmological gravitational particle production (GPP) Parker:1969au ; Parker:1971pt ; Ford:2021syk ; Chung:1998zb ; Chung:1998ua ,and arises from the dynamics of a quantum field on a curved spacetime background DeWitt:1975ys ; Birrell:1982ix ; Parker:2009uva . GPP is a generic process with very few necessary ingredients, only requiring a massive spectator field on top of an inflating spacetime background. GPP is particularly relevant in the context of dark matter, which is known to make up a quarter of the energy density of the universe, though its exact particle nature is unknown Planck:2018jri . As it requires no coupling to standard model particles, GPP is an attractive candidate to produce some or all of the dark matter, yielding ‘completely dark’ matter. GPP has been studied for a wide variety of inflation models and for particles ranging from scalars to spin-s, e.g., Chung:1998zb ; Graham:2015rva ; Ema:2015dka ; Ema:2016hlw ; Ema:2018ucl ; Ema:2019yrd ; Kolb:2020fwh ; Ahmed:2020fhc ; Alexander:2020gmv ; Kolb:2021xfn ; Kolb:2021nob ; Kolb:2022eyn ; Kolb:2023dzp ; Capanelli:2023uwv ; Kaneta:2023uwi ; Maleknejad:2022gyf ; Capanelli:2024pzd ; Capanelli:2024nkf ; Racco:2024aac ; Choi:2024bdn ; Verner:2024agh (see Kolb:2023ydq and references therein for a more detailed review). For each of these models there is a region of parameter space in which particles produced gravitationally can be the dark matter, but GPP can also be a relevant production mechanism for other cosmological relics.
A variety of methods have been employed to study GPP. In some cases, it is possible to approximate particle production using various analytical methods, but for a full analysis, numerical methods must be employed. In order to understand the spectra of produced particles, it is thus important to fully understand where each analytic method is valid and how these approximations compare to the full numerical calculation. The relevant analytic methods are as follows: first, one can trace out the evolution of each mode as it transitions from the inflationary quasi-de Sitter regime into matter and/or radiation domination by solving the mode equation in each region and matching at the boundaries. This method we refer to as ‘boundary matching.’ One can also employ the Stokes phenomenon Hashiba:2021npn and steepest descent PhysRevD.67.083514 methods to approximate GPP. Finally, GPP can be analytically approximated by considering inflaton scattering processes, as discussed in e.g. Tang_2017 ; Basso_2022 . In this work, we perform a detailed analysis of the GPP for scalars, considering both analytic and numerical methods for the GPP of both conformally and minimally-coupled scalars. For the background inflationary dynamics we consider quadratic inflation and both early and late reheating scenarios. We analytically approximate the GPP in each of these cases and outline the regime of validity of each analytic method. We additionally numerically calculate the GPP and compare with the analytic results, finding good agreement between the two.
The structure of the paper is as follows: in Sec. 2 we provide a brief overview of the GPP procedure and the setup for a massive scalar field on an inflationary background. In Sec. 3, we present comprehensive numerical results for the GPP of scalars, considering both conformal and minimal couplings as well as both early and late reheating. We then turn to a comparison with analytic results, summarizing the analytic methods and results in Sec. 4. Finally we conclude with a discussion in Sec. 5. Technical details of the analytic calculations can be found in the appendices.
Throughout the paper, the following conventions are used: we use a mostly minus metric signature and employ natural units such that $c=\hbar=1$. Greek indices, $\mu,\nu,\ldots$ indicate a sum over all four spacetime indices, while Latin indices indicate a spatial sum only. We furthermore take ${M_{\mathrm{Pl}}}=(8\pi G)^{-1/2}$.
2 Inflation and Gravitational Particle Production
In this section we provide a brief overview of the mechanics of GPP and the assumptions of the background evolution of the Universe during inflation. Detailed descriptions of GPP be found in several textbooks, such as Birrell:1982ix ; Mukhanov:2005sc ; Mukhanov:2007zz ; Parker:2009uva ; Baumann:2022mni , as well as more recent reviews, such as Kolb:2023ydq ; Ford:2021syk ; Armendariz-Picon:2023gyl , so here we focus just on the key features.
GPP is a semiclassical process, meaning that the background gravitational field is classical, whereas the spectator field is quantized. For the background evolution, we assume a cosmological spacetime that is homogeneous, isotropic, and expanding. This background is described by the spatially flat Friedmann-Robertson-Walker (FRW) metric with line element given by $ds^{2}=a^{2}(\eta)\left(d\eta^{2}-d{\bm{x}}^{2}\right)$, where $\eta$ denotes conformal time, $a$ denotes scale factor, and the two quantities are related by $a\,d\eta=dt$. For the early evolution of the Universe, we assume starting in a (quasi-)de Sitter (dS) phase of accelerated expansion (inflation), followed by a period of evolution as a matter-dominated (MD) universe, then “reheating” to a radiation-dominated (RD) Universe.
This evolution is conveniently engineered by assuming that there is a single scalar field - the inflaton field $\varphi$, whose potential energy drives an inflationary phase with $\ddot{a}>0$ (dot is derivative with respect to coordinate time). Inflation ends when the kinetic energy of the inflaton field dominates the potential energy and $\ddot{a}$ first becomes negative. The inflaton field then undergoes coherent oscillations about the minimum of its potential, and its energy density redshifts as matter ($\rho_{\varphi}\propto a^{-2/3}$) when averaged over several oscillations. During this MD era, the energy density evolves as $\exp{(-\Gamma_{\varphi}t)}$, where the decay width, $\Gamma_{\varphi}$, is the rate at which the inflaton decays into relativistic standard model particles. Eventually, the radiation dominates the inflaton energy density, at which point the Universe “reheats” and enters the RD epoch.
We consider two reheating scenarios: early reheating and late reheating. Late reheating occurs when $(a_{\mbox{\tiny RH}}/a_{e})>(m/H_{e})^{-2/3}$, i.e., when $H_{\mbox{\tiny RH}}<m$. In numerical calculations, late reheating is implemented by setting $\Gamma_{\varphi}=0$. For early reheating, there exists a limiting case of “immediate reheating,” where reheating occurs instantaneously at the end of inflation.
For reheating we follow the analysis in Ref. PhysRevD.64.023508 . Although the conversion of the inflaton energy density into radiation is effective throughout the MD and RD eras, we use “reheating” to refer to the moment where $\rho_{\varphi}=\rho_{\mathrm{R}}$. The radiation energy density at reheating can be expressed as $\rho_{\mbox{\tiny RH}}=(\pi^{2}g_{\star}/30)T_{\mbox{\tiny RH}}^{4}$, where $g_{\star}$ is the effective number of relativistic degrees of freedom at $T_{\mbox{\tiny RH}}$. From the Friedmann equation, we obtain $3{M_{\mathrm{Pl}}}^{2}H_{\mbox{\tiny RH}}^{2}=\rho_{\mathrm{R}}(T_{\mbox{\tiny RH}})$. By setting $\rho_{\varphi}=\rho_{R}$ and $H=\Gamma_{\varphi}$, the reheat temperature can be expressed as
$T_{\mbox{\tiny RH}}=\left(\Gamma_{\varphi}{M_{\mathrm{Pl}}}\right)^{1/2}\left(\frac{90}{\pi^{2}g_{\star}}\right)^{1/4}\ .$ (1)
The production of neutrinos during nucleosynthesis places a lower bound of $T_{{\mbox{\tiny RH}}}^{\rm min}\approx 5$ de_Salas_2015 . The maximum $T_{{\mbox{\tiny RH}}}$ depends the energy scale of inflation, as
$T_{\mbox{\tiny RH}}<\left(H_{e}{M_{\mathrm{Pl}}}\right)^{1/2}\left(\frac{90}{\pi^{2}g_{\star}}\right)^{1/4}$ (2)
For inflation itself, there are a wide range of models describing the inflationary dynamics (see, e.g., martin2023encyclopaedia for a summary and review). For simplicity, we here we adopt a simple model of single-field quadratic inflation. The action for the inflaton field in the quadratic model is
$S_{\varphi}=\int\mathrm{d}^{4}x\sqrt{-g}\left[\tfrac{1}{2}g^{\mu\nu}\nabla_{\mu}\varphi\nabla_{\nu}\varphi-\tfrac{1}{2}m_{\varphi}^{2}\varphi^{2}\right]\ .$ (3)
The energy density $\rho_{\varphi}$ and pressure $p_{\varphi}$ are given by
$\rho_{\varphi}=\frac{1}{2}\dot{\varphi}^{2}+\frac{1}{2}m_{\varphi}^{2}\varphi^{2}\ ;\quad p_{\varphi}=\frac{1}{2}\dot{\varphi}^{2}-\tfrac{1}{2}m_{\varphi}^{2}\varphi^{2}\ ,$ (4)
and the equation of motion for the inflaton field is
$\ddot{\varphi}+3H\dot{\varphi}+m_{\varphi}^{2}\varphi+\Gamma_{\varphi}\dot{\varphi}=0$ (5)
The Hubble expansion rate is $H=\dot{a}/a=a^{\prime}/a^{2}$ (prime denotes derivative with respect to conformal time), and the Ricci curvature is $R=-6\ddot{a}/a-6\dot{a}^{2}/a^{2}=-6a^{\prime\prime}/a^{3}$. As discussed above, the decay width of the inflaton, $\Gamma_{\varphi}$, sets the dynamics of reheating.
We use the subscript “$e$” to denote the end of inflation, and the subscript “RH” to denote reheating. For example, $a_{\mbox{\tiny RH}}$ denotes the value of the scale factor at reheating, and $a_{e}$ denotes the value of the scale factor at the end inflation. The dS phase lasts between $\eta\rightarrow-\infty$ and $\eta_{e}$, the MD phases lasts between $\eta_{e}$ and $\eta_{\mbox{\tiny RH}}$, and the RD phase commences at $\eta_{\mbox{\tiny RH}}$.
For convenience, we will often rescale common variables relative to $a_{e}$, $H_{e}$, and $\eta_{e}$, which respectively denote the values at the end of inflation of the scale factor, Hubble constant, and conformal time. We define the following dimensionless quantities: ${\tilde{a}}=a/a_{e}$, ${\tilde{m}}=m/H_{e}$, ${\tilde{k}}=k/a_{e}H_{e}$, ${\tilde{\eta}}=a_{e}H_{e}(\eta-\eta_{e})-1$, ${\tilde{\omega}}_{k}=\omega_{k}/a_{e}H_{e}$, ${\tilde{H}}=H/H_{e}$, ${\tilde{n}}_{k}=n_{k}/H_{e}^{3},{\tilde{n}}=n/H_{e}^{3}$, and $\tilde{R}=R/6H_{e}^{2}$.
For our analytic calculations, we will often approximate the evolution of the background spacetime. In Table 2 we summarize how the scale factor $\tilde{a}$, the scalar curvature $\tilde{R}$, and the Hubble parameter $\tilde{H}$ depend on the conformal time $\tilde{\eta}$ in de Sitter evolution, MD evolution, and RD evolution. Our conventions are such that at the end of inflation, $\tilde{\eta}_{e}=-1$, and $\tilde{a}_{e}=1$.
de Sitter (dS) Matter domination (MD) Radiation domination (RD)
$\tilde{a}$ $-\tilde{\eta}^{-1}$ $\left(\dfrac{3}{2}+\dfrac{\tilde{\eta}}{2}\right)^{2}$ ${\tilde{a}}_{\mbox{\tiny RH}}+{\tilde{a}}_{\mbox{\tiny RH}}^{1/2}({\tilde{\eta}}-{\tilde{\eta}}_{\mbox{\tiny RH}})\approx{\tilde{a}}_{\mbox{\tiny RH}}^{1/2}{\tilde{\eta}}\quad({\tilde{\eta}}\gg{\tilde{\eta}}_{\mbox{\tiny RH}})$
$\tilde{R}$ $-2$ $-\dfrac{1}{2}\tilde{a}^{-3}=-\dfrac{1}{2}\left(\dfrac{3}{2}+\dfrac{\tilde{\eta}}{2}\right)^{-6}$ 00
$\tilde{H}$ $1$ $\tilde{a}^{-3/2}=\left(\dfrac{3}{2}+\dfrac{\tilde{\eta}}{2}\right)^{-3}$ $\tilde{a}_{\mbox{\tiny RH}}^{1/2}\tilde{a}^{-2}\approx\tilde{a}_{\mbox{\tiny RH}}^{-1/2}\tilde{\eta}^{-2}\quad({\tilde{\eta}}\gg{\tilde{\eta}}_{\mbox{\tiny RH}})$
[…]
tableDependence of scale factor ${\tilde{a}}$, Hubble constant ${\tilde{H}}$, and scalar curvature $\tilde{R}$ on conformal time ${\tilde{\eta}}$ during the epochs of quasi de Sitter (dS), matter domination (MD), and radiation domination (RD).
In addition to the inflaton field determining the evolution of the background spacetime, we consider an additional “spectator” scalar field $\sigma$ which does not effect the evolution of the background. The action is
$S_{\sigma}=\int\mathrm{d}^{4}x\sqrt{-g}\left[\tfrac{1}{2}g^{\mu\nu}\nabla_{\mu}\sigma\nabla_{\nu}\sigma-\tfrac{1}{2}m_{\sigma}^{2}\sigma^{2}+\tfrac{1}{2}\xi R\sigma^{2}\right]\ .$ (6)
We will consider two types of coupling to gravity: minimal, where the coupling constant $\xi=0$, and conformal, where the coupling constant $\xi=1/6$.
In an FRW spacetime the kinetic term for $\sigma$ is noncanonical, so one performs a change of variables $\sigma(\eta,{\bm{x}})=a^{-1}(\eta)\phi(\eta,{\bm{x}})$ and the action becomes
$S^{\mathrm{FRW}}[\phi(\eta,{\bm{x}})]$ $=\int\!\mathrm{d}\eta\int\!\mathrm{d}^{3}{\bm{x}}\,\Bigl{[}\tfrac{1}{2}(\phi^{\prime})^{2}-\tfrac{1}{2}|\bm{\nabla}\phi|^{2}-\tfrac{1}{2}a^{2}m_{\mathrm{eff}}^{2}\phi^{2}\Bigr{]}\,,$ (7)
$m_{\mathrm{eff}}^{2}(\eta)$ $=m^{2}+(\tfrac{1}{6}-\xi)R(\eta)\ ,$ (8)
where $\bm{\nabla}=(\partial_{x},\partial_{y},\partial_{z})$ denotes derivatives with respect to comoving spatial coordinates. Note that the effective mass squared $a^{2}(\eta)m_{\mathrm{eff}}^{2}(\eta)$ is time-dependent and the sign can be positive or negative. In conformal time, the equation of motion is
$\phi^{\prime\prime}-\bm{\nabla}^{2}\phi+a^{2}m_{\mathrm{eff}}^{2}\phi=0\ ,$ (9)
which is of the form of a second-order linear wave equation with a time-dependent effective mass.
Next, we promote $\phi$ to a quantum operator with canonical commutation relations, and write the solution to the linear wave equation in terms of mode functions $\chi_{\bm{k}}(\eta)$:
$\hat{\phi}(\eta,{\bm{x}})=\int\!\!\frac{\mathrm{d}^{3}{\bm{k}}}{(2\pi)^{3}}\Bigl{[}\hat{a}_{\bm{k}}\,\chi_{\bm{k}}(\eta)\,e^{\mathrm{i}{\bm{k}}\cdot{\bm{x}}}+\hat{a}_{\bm{k}}^{\dagger}\,\chi_{\bm{k}}^{*}(\eta)\,e^{-\mathrm{i}{\bm{k}}\cdot{\bm{x}}}\Bigr{]}\ .$ (10)
The ladder operators $\hat{a}_{\bm{k}}$ and $\hat{a}_{\bm{k}}^{\dagger}$ satisfy the usual commutation relations $[\hat{a}_{\bm{k}},\hat{a}^{\dagger}_{\bm{q}}]=(2\pi)^{3}\delta({\bm{k}}-{\bm{q}})$, $[\hat{a}_{\bm{k}},\hat{a}_{\bm{q}}]=[\hat{a}_{\bm{k}}^{\dagger},\hat{a}_{\bm{q}}^{\dagger}]=0$ The mode equation becomes [^1] Now we take advantage of the isotropy of the background spacetime and note that we need only consider $k\equiv|{\bm{k}}|$.
$\chi_{k}^{\prime\prime}(\eta)+\omega_{k}^{2}(\eta)\chi_{k}(\eta)=0\ .$ (11)
This is a harmonic oscillator equation for each wavenumber $k$ with a time dependent frequency given by
$\omega_{k}^{2}(\eta)=k^{2}+a^{2}(\eta)m^{2}+\bigl{(}\tfrac{1}{6}-\xi\bigr{)}a^{2}(\eta)R(\eta)\ .$ (12)
Gravitational particle production arises if the initial (early) and final (late) [^2] Note that in this context, ‘early’ and ‘late’ refer to the initial and final states and are distinct from early and late reheating. creation and annihilation operators are not the same. We define the initial vacuum state with early-time creation and annihilation operators $\hat{a}_{k}^{{\mbox{\tiny early}}\dagger}$ and $\hat{a}_{k}^{{\mbox{\tiny early}}}$ by
$\hat{a}_{k}^{\mbox{\tiny early}}\ket{0^{\mbox{\tiny early}}}=0\ket{0^{\mbox{\tiny early}}}\ .$ (13)
It is related to the late-time operators by
$\hat{a}_{k}^{\mbox{\tiny early}}=\alpha_{k}^{*}\hat{a}_{k}^{\mbox{\tiny late}}-\beta_{k}^{*}\hat{a}_{-k}^{{\mbox{\tiny late}}\dagger},$ (14)
where $\alpha_{k}$ and $\beta_{k}$ are Bogoliubov coefficients. The value of $\beta_{k}$ can be found from the mode equations such that
$|\beta_{k}|^{2}=\lim_{\eta\rightarrow\infty}\left[\frac{\omega_{k}}{2}|\chi_{k}|^{2}+\frac{1}{2\omega_{k}}|\partial_{\eta}\chi_{k}|^{2}-\frac{1}{2}\right]\ ,$ (15)
where $\omega_{k}$ is the frequency of the mode and the factor of 1/2 arises from the normalization of the field. Note that the expression Eq. (15) is only valid for $\omega_{k}^{2}>0$. From $|\beta_{k}|^{2}$ we can construct the spectrum of particles produced Kolb:2020fwh
$n_{k}=\frac{k^{3}}{2\pi^{2}}|\beta_{k}|^{2}\ ,$ (16)
where $k$ is the comoving wavenumber. Then, the comoving number density, $a^{3}n$ is given by
$a^{3}n=\int\frac{dk}{k}n_{k}\ .$ (17)
From the comoving number density, one can determine the relic density.
We will take the Bunch-Davies initial condition to solve for particle production. In the early time limit $a\rightarrow 0$ and $a^{2}R\rightarrow$0, which implies that $\omega_{k}^{2}\rightarrow k^{2}$. In this limit, the mode equation has solution
$\lim_{k\eta\rightarrow-\infty}\chi_{k}(\eta)=\frac{1}{\sqrt{2k}}e^{-ik\eta}\ .$ (18)
3 Numerical Results
In this section we present our numerical GPP results. We first describe the distinctions between the early and late reheating scenarios, then present results for conformally-coupled scalars in Sec. 3.1 and minimally-coupled scalars in Sec. 3.2.
In Fig. 1, we compare the evolution of ${\tilde{H}}$ for three different values of $\Gamma_{\varphi}$, which correspond to late reheating ($\Gamma_{\varphi}=0$) and two nonzero values for $\Gamma_{\varphi}$, corresponding to early reheating. We emphasize that though we have defined ‘late reheating’ to correspond to $\Gamma_{\varphi}=0$, in reality it corresponds to a range of parameters such that ${\tilde{H}}_{\mbox{\tiny RH}}<m$. One can appreciate this from Fig. 1, in which the reheating scenarios diverge from each other at ${\tilde{a}}={\tilde{a}}_{\mbox{\tiny RH}}$.
Figure 1: Comparison of $H/H_{e}$ for late (blue solid line) and early (red and cyan dashed lines) reheating.
The difference between early and late reheating is largely driven by the difference in the late time evolution of $H$. A later $\tilde{a}_{\mbox{\tiny RH}}$ therefore corresponds to an early reheating background that is closer to that of late reheating. For the remainder of this work, we take $\tilde{a}_{\mbox{\tiny RH}}=3.34$, corresponding to a decay width of $\Gamma_{\varphi}/m_{\varphi}=0.1$, for our early reheating numerical analysis. For the analytic analysis, $a_{\mbox{\tiny RH}}$ will be a free parameter.
To determine the evolution of $\tilde{a}^{3}\tilde{n}_{k}$, we solve the differential equation for $\chi_{k}$ and $\chi_{k}^{\prime}$ in $\eta$, starting from deep in the relativistic region ($\tilde{k}>\tilde{a}\tilde{m}$) with Bunch-Davies initial conditions and ending where the oscillations in the Bogoliubov coefficient (and thus the value of $\tilde{a}^{3}\tilde{n}_{k}$) have “frozen out” to a constant value, which is used as the final value for $\tilde{a}^{3}\tilde{n}_{k}$.
Example numerical results of the evolution of ${\tilde{a}}^{3}{\tilde{n}}_{k}$ in ${\tilde{a}}$ are shown in Fig. 2 for ${\tilde{m}}=0.003$ and two choices of ${\tilde{k}}=0.03$ and ${\tilde{k}}=0.3$ for both late and early reheating. In both cases conformal coupling ($\xi=1/6$) was assumed. The left figure is ${\tilde{a}^{3}\tilde{n}_{k}}$ versus $\tilde{a}$ for $\tilde{k}=0.03$ (so $\tilde{k}<\tilde{m}^{1/3}$), and the right figure is ${\tilde{a}^{3}\tilde{n}_{k}}$ versus ${\tilde{a}}$ for ${\tilde{k}}=0.3$ (so $\tilde{k}>\tilde{m}^{1/3}$). For $\tilde{m}\geq 1$, the spectra would be the same for early and late reheating.
Figure 2: Comparison of the evolution of $\tilde{a}^{3}\tilde{n}_{k}$ in $\tilde{a}$ for late and early reheating, for conformally coupled scalars with $\tilde{m}=0.003$ and the indicated values of $\tilde{k}$.
For minimally-coupled scalars, the calculation is slightly more subtle, because $\omega_{k}$ (Eq. (12)), which enters Eq. 15, can be imaginary if the $R$ term dominates. Because Eq. (15) is only valid for $\omega_{k}^{2}>0$, we emphasize that $\tilde{a}^{3}\tilde{n}_{k}$ only has physical meaning at late times (see Kolb:2023ydq for more discussion). However, it still exhibits the same “freeze out” behavior, and the same numerical procedure described above is used to determine $\tilde{a}^{3}\tilde{n}_{k}$ values for minimally-coupled scalars.
3.1 Conformal coupling
Let us now discuss the features of the spectrum in detail. We first consider conformally-coupled scalars. Shown in Fig. 3 is an example of the $\tilde{a}^{3}\tilde{n}_{k}$ versus $\tilde{k}$ spectra for conformally-coupled scalars with masses ranging from $\tilde{m}=10^{-4}$ to $\tilde{m}=10^{-1}$. Late and early reheating are denoted by solid and dotted lines, respectively. As expected, the early and late results approach each other as $\tilde{m}$ approaches $1$. We also show $\tilde{a}^{3}\tilde{n}_{k}$ evaluated at the “dominant modes” ${\tilde{k}}_{\star}$ in orange, which approximate the peak values of the spectra. We will discuss this further in Sec. 3.1.1.
Figure 3: Numerical $\tilde{a}^{3}\tilde{n}_{k}$ versus $\tilde{k}$ spectra of conformally-coupled scalars. Solid and dotted lines represent late and early reheating, respectively. Note that at low ${\tilde{k}}$, late reheating spectra are dominant over those of early reheating, but become subdominant in the red-tilted region. The orange curves trace the value of the spectra evaluated at the “dominant modes” ${\tilde{k}}_{\star}$, which approximate the peak values of the spectra.
After obtaining $\tilde{a}^{3}\tilde{n}_{k}$, we can find the comoving number density, $\tilde{a}^{3}\tilde{n}$. For completeness, one should perform the full integral over $\tilde{n}_{k}$ as in Eq. (17). However, for conformally-coupled scalars, one can simply approximate the full integral as the peak value of $\tilde{a}^{3}\tilde{n}_{k}$, which we denote as $\left[{\tilde{a}^{3}\tilde{n}_{k}}\right]_{\mathrm{peak}}$.
Figure 4: Comoving number density ($\tilde{a}^{3}\tilde{n}$) for a range of masses, with analytic fits that approximate the data points. Late reheating (red) and early reheating (blue) values agree for $\tilde{m}\geq 0.1$. The dotted lines represent ${\tilde{a}^{3}\tilde{n}_{k}}$ evaluated at ${\tilde{k}}_{\star}$, which is discussed in Sec. 3.1.1.
Figure 4 shows the comoving number density $\tilde{a}^{3}\tilde{n}$ for a range of masses from $\tilde{m}=10^{-4}$ to $\tilde{m}=10$. We show the numerical results for late and early reheating values as red dots and blue dots, respectively, as well as the best fit lines to the numerical results. For small masses ($\tilde{m}\leq 1$), the comoving number density for late (early) reheating scales as $\tilde{m}$ ($\tilde{m}^{3/2}$), until $\tilde{m}\approx 0.1$, where the results for late and early reheating converge. For $\tilde{m}\geq 1$, the comoving number density for both late and early reheating scale as $e^{-2.1\tilde{m}}$. For each mass, we also plot the value of ${\tilde{a}^{3}\tilde{n}_{k}}$ evaluated at the dominant mode, $k_{\star}$. We see that the numerically integrated value of ${\tilde{a}}^{3}{\tilde{n}}$ is well approximated (up to a factor of two) by assuming $\tilde{a}^{3}\tilde{n}\approx\left[{\tilde{a}^{3}\tilde{n}_{k}}\right]_{{\tilde{k}}_{\star}}$.
3.1.1 Dominant modes
Conformally-coupled scalars have spectra that are strongly peaked, so that $\tilde{a}^{3}\tilde{n}$ can be approximated by the value of $\tilde{a}^{3}\tilde{n}_{k}$ at $\tilde{k}_{\mathrm{peak}}$. These are the modes that carry most of the comoving number density, and can in turn be approximated by the “dominant mode” of the spectrum, which we denote by $\tilde{k}_{\star}$. Dominant modes exist for conformally-coupled light mass scalars ($\tilde{m}<1$ and $\xi=1/6$). These are modes that re-enter the horizon after inflation at ${\tilde{a}}={\tilde{a}}_{\star}$ when $3\tilde{H}({\tilde{a}}_{\star})=\tilde{m}$, and can be found as ${\tilde{k}}_{\star}={\tilde{a}}_{\star}{\tilde{m}}/3$.
During matter domination, we approximate $\tilde{H}({\tilde{a}}_{\star})=\tilde{a}_{\star}^{-3/2}$, yielding ${\tilde{a}}_{\star}=({\tilde{m}}/3)^{-2/3}$ and ${\tilde{k}}_{\star}=({\tilde{m}}/3)^{1/3}$. During radiation domination, we approximate $\tilde{H}_{\star}=\tilde{a}_{\mbox{\tiny RH}}^{1/2}\tilde{a}_{\star}^{-2}$, yielding ${\tilde{a}}_{\star}=(3{\tilde{a}}_{\mbox{\tiny RH}}^{1/2}/{\tilde{m}})^{1/2}$ and ${\tilde{k}}_{\star}={\tilde{a}}_{\mbox{\tiny RH}}^{1/4}({\tilde{m}}/3)^{1/2}$.
Figure 5 compares the numerically determined peak value in ${\tilde{k}}$ of $\tilde{a}^{3}\tilde{n}_{k}$ and the value of the dominant mode $\tilde{k}_{\star}$ for four values of $\tilde{m}$ from $\tilde{m}=10^{-4}$ to $\tilde{m}=10^{-1}$. Indicated in blue are modes that re-enter horizon during matter domination, and represented by red are the results for early reheating spectra. Up to order unity, there is agreement between the peaks found via the numerical spectra and the dominant modes derived above. Finally, we note that we will usually drop the factors of $3$ in the definitions of ${\tilde{k}}_{\star}$, and refer to ${\tilde{k}}_{\star}={\tilde{m}}^{1/3}$ and ${\tilde{k}}_{\star}={\tilde{a}}_{\mbox{\tiny RH}}^{1/4}{\tilde{m}}^{1/2}$ as the dominant modes for late and early reheating respectively.
Figure 5: Comparison the the results for the dominant mode (${\tilde{k}}_{\star}$–solid curves) and numerical results for the peak value (points) for late (blue) and early (red) reheating scenarios.
3.2 Minimal coupling
Having enumerated the features of the spectrum for conformally-coupled scalars, let us now discuss minimally-coupled scalars. In Fig. 6 we present numerical results for $\tilde{a}^{3}\tilde{n}_{k}$ as a function of $\tilde{k}$ for minimally-coupled scalars with masses $\tilde{m}=10^{-4}$ to $\tilde{m}=10^{-1}$ for late (solid) and early (dotted) reheating. As we noted before for conformally-coupled scalars, the early and late reheating curves approach each other as $\tilde{m}$ approaches $1$. The $\tilde{k}$ range is chosen from $10^{-4}$ to $1$, so that we can see that the spectra tend toward flatness at low $\tilde{k}$, and fully converge around $\tilde{k}=1$.
Figure 6: Spectra of minimally-coupled scalars, late (solid) and early (dotted) reheating.
Figure 7: Comoving number density ($\tilde{a}^{3}\tilde{n}$) for a range of masses, with approximate fits to the numerical points. Late reheating (red) and early reheating (blue) agree for $\tilde{m}\geq 0.5$.
We now again find the comoving number density $\tilde{a}^{3}\tilde{n}$ from $\tilde{a}^{3}\tilde{n}_{k}$. Unlike for conformally-coupled scalars where the spectra are blue-tilted before reaching ${\tilde{k}}_{\mathrm{peak}}$, the spectra of minimally-coupled scalars are red-tilted in the low-${\tilde{k}}$ region, and continue to decrease as ${\tilde{k}}$ increases. This means that the $\tilde{k}$ range we have calculated numerically only captures the “tail end” of the area under the curve, and modes with smaller ${\tilde{k}}$ carry more of the number density. Therefore, before finding $\tilde{a}^{3}\tilde{n}$ by integrating the ${\tilde{a}^{3}\tilde{n}_{k}}$ spectrum, we first extrapolate the spectrum to ${\tilde{k}}=10^{-20}$ by fitting a line to the low-${\tilde{k}}$ region. This value of minimum ${\tilde{k}}$ is chosen since CMB observations probe scales of approximately $10^{4}{\tilde{k}}_{0}$, where ${\tilde{k}}_{0}=a_{0}H_{0}/a_{e}H_{e}$ are the modes that are of a scale equal to the Hubble scale today. Using $a_{0}/a_{\mbox{\tiny RH}}=T_{\mbox{\tiny RH}}/T_{0}$ and $a_{\mbox{\tiny RH}}/a_{e}=(H_{e}/H_{\mbox{\tiny RH}})^{3/2}$, we find ${\tilde{k}}_{0}\sim 10^{-24}(10^{9}{\mathrm{GeV}}/T_{\mbox{\tiny RH}})^{1/3}(10^{12}{\mathrm{GeV}}/H_{e})^{1/3}$. Therefore, the spectrum is extended to ${\tilde{k}}=10^{4}{\tilde{k}}_{0}=10^{-20}$.
Figure 7 shows $\tilde{a}^{3}\tilde{n}$ for masses ranging from $10^{-4}$ to $3$. For $\tilde{m}<1$, $\tilde{a}^{3}\tilde{n}$ decreases as $\tilde{m}^{-1}$ for late reheating, and as ${\tilde{m}}^{-1/2}$ for early reheating. For $\tilde{m}>1$, $\tilde{a}^{3}\tilde{n}$ exponentially decreases in $\tilde{m}$. After $\tilde{m}\approx 0.5$, the late reheating and early reheating $\tilde{a}^{3}\tilde{n}$ agree. As we will discuss further in the next section, the analytic and numerical results agree overall.
4 Analytic Methods and Results
Having numerically calculated the GPP, we now turn to the analytic approximations. We first provide a brief explanation of the various analytic methods and their regimes of validity in Sec. 4.1. We then present our results for conformally and minimally-coupled scalars in Sec. 4.2 and Sec. 4.3, respectively, as well as a comparison between the analytic approximations and numerical results.
4.1 Description of Analytic Methods
We now outline the various methods available to analytically approximate particle production: boundary matching, Stokes phenomenon, steepest descent, and inflaton scattering. Here we provide a brief summary and discuss where each method is valid. We will be particularly interested in the $\tilde{k}$ and $\tilde{m}$ scaling of the solutions. Further details of each method are outlined in the Appendices and relevant references.
  1. 1.
    Boundary Matching: In the boundary matching method, we find solutions to the mode equation as it evolves through the quasi-de Sitter phase into radiation and/or matter domination, such that the boundary conditions are set by the solution in the previous epoch. The modes and their derivatives are then matched at each boundary. This method is applicable when $\tilde{k}\ll 1$ and $\tilde{m}\ll 1$, and detailed calculations are found in Appendix A.
  2. 2.
    Stokes Phenomenon: The mixing of mode functions leading to particle production can also be understood in terms of the Stokes phenomenon. We follow Ref. Hashiba:2021npn in applying the Stokes phenomenon to GPP. The Stokes phenomenon is applicable when ${\tilde{k}}>{\tilde{m}}^{1/3}$ and ${\tilde{k}}>{\tilde{m}}$. See Appendix B for the calculations. The Stokes approach will only be used for conformal coupling.
  3. 3.
    Steepest Descent: Similar to the Stokes phenomenon, the steepest descent method approximates particle production by evaluating the integral for $\beta_{k}$ around its poles in the complex ${\tilde{\eta}}$ plane. This method is described in Ref. PhysRevD.67.083514 . The steepest descent method can be applied when $\tilde{k}>\tilde{m}^{1/3}$ for light particles ($\tilde{m}<1$), where it agrees with the results of the Stokes phenomenon method, and for $\tilde{k}<\tilde{m}$ for heavy particles ($\tilde{m}>1$). See Appendix C for the calculations.
  4. 4.
    Inflaton Scattering: As the name suggests, this method approximates particle production as a result of inflaton annihilation into two $\chi$ particles. For particles with $m\leq m_{\varphi}$, the dominant channel is $\varphi\varphi\rightarrow\chi\chi$, while for particles with $m>m_{\varphi}$, one must consider $n\varphi\rightarrow\chi\chi$, where $n$ is the smallest even number such that $m<nm_{\varphi}/2$. This process is kinematically allowed when $m<m_{\varphi}$, and is the dominant production mechanism when $k\gtrsim m_{\varphi}$. For quadratic inflation, $m_{\varphi}\approx 2H_{e}$, while for Hilltop inflation with potential $V(\varphi)=(m_{\varphi}^{2}v^{2}/72)(1-\varphi^{6}/v^{6})^{2}$, $m_{\varphi}\approx 31H_{e}$ Basso_2022 . For the rest of this study, we consider the quadratic inflation model. We find numerically that inflaton scattering becomes effective at ${\tilde{k}}\gtrsim 1$ for ${\tilde{m}}\lesssim 1$, and at ${\tilde{k}}\gtrsim 2$ for ${\tilde{m}}\gtrsim 1$. See Appendix D for the calculations and Refs. Tang_2017 ; Basso_2022 for further discussions.
4.2 Analytic Results: Conformal Coupling
We now discuss the application of the above analytic approximations to calculating ${\tilde{a}}^{3}{\tilde{n}}_{k}$ for conformally coupled scalars for both early and late reheating scenarios. It is convenient to characterize the particle production in terms of the $\tilde{m}-\tilde{k}$ plane, shown in Fig. 8 for late (left) and early (right) reheating. For late reheating, the particle production can be divided into regions delineated by ${\tilde{k}}_{\star}={\tilde{m}}^{1/3}$ and ${\tilde{k}}={\tilde{m}}$, while for early reheating, the regions are delineated by ${\tilde{k}}={\tilde{m}}$, ${\tilde{k}}=\tilde{a}_{\mbox{\tiny RH}}{\tilde{m}}$, and ${\tilde{k}}_{\star}=\tilde{a}_{\mbox{\tiny RH}}^{1/4}{\tilde{m}}^{1/2}$.
Figure 8: Regions of ${\tilde{k}}$ vs. ${\tilde{m}}$ for conformally-coupled scalars for late reheating (left) and early reheating (right). The methods for each region of the left panel are as follows: Region O: inflaton scattering, Region I: steepest descent and Stokes phenomenon, Regions II-III: boundary matching, and Region V: steepest descent. Region IV is a transition region between the behaviors of Regions O and V. The methods for each region of the right panel are as follows - Region O: inflaton scattering, Region I: Stokes phenomenon, and Regions II-IV: steepest descent.
For each of the numbered regions we analytically approximate the particle production. Tables 4.2 and 4.2 show the results ${\tilde{a}}^{3}\tilde{n}_{k}$ for late and early reheating, respectively. We show the results of each method in the regions where it is applicable, as well as the best fit to the numerical results found in Sec. 3. We focus primarily on the scaling of ${\tilde{a}}^{3}{\tilde{n}}_{k}$ with ${\tilde{k}}$ and ${\tilde{m}}$ and collect numerical factors into constants. Each table is further divided into ${\tilde{m}}<1$ and ${\tilde{m}}>1$.
For late reheating, summarized in Table 4.2, we see that for ${\tilde{m}}<1$, $2\rightarrow 2$ scattering is relevant in Region O, both the Stokes phenomenon method and steepest descent can be applied in Region I, and boundary matching in Regions II and III. In each of these cases, the analytic results reproduce the numerical scaling in ${\tilde{k}}$ and ${\tilde{m}}$ up to numerical factors. For ${\tilde{m}}>1$, we now have $n\rightarrow 2$ scattering for ${\tilde{m}}<n$ (e.g. $4\rightarrow 2$ for $42<{\tilde{m}}<$ 4) in Region O, and the steepest descent method applies in Region V. In this scenario, Region IV only exists from $21<{\tilde{k}}<$ 2 and $21<{\tilde{m}}<$ 2, and the numerical results show that spectrum of Region IV appears as a transitional region between the $\sim{\tilde{k}}^{3}$ behavior of Region V and the $\sim{\tilde{k}}^{-3/2}$ behavior of Region O.
The early reheating results are shown in Table 4.2. In this scenario, we only consider small mass, ${\tilde{m}}<1$. We see that in Regions II-IV, the ${\tilde{k}}^{2}$ scaling of the numerical results is well captured by the boundary matching method. In Region O, we numerically find an exponential suppression, which is well approximated by $2\rightarrow 2$ scattering. In this case, we individually consider the contribution to scattering during the MD and RD eras, demarcated by ${\tilde{k}}={\tilde{m}}_{\varphi}{\tilde{a}}_{\mbox{\tiny RH}}$. We define $\Phi_{\mbox{\tiny MD}}$ and $\Phi_{\mbox{\tiny RD}}$:
$\Phi_{\mbox{\tiny MD}}$ $=\exp\left(-\dfrac{2\Gamma_{\varphi}\tilde{k}^{3/2}}{3H_{e}\tilde{m}_{\varphi}^{3/2}}\right)$ $({\tilde{k}}<{\tilde{m}}_{\varphi}{\tilde{a}}_{\mbox{\tiny RH}})\ ,$
$\Phi_{\mbox{\tiny RD}}$ $=\exp\left(-\dfrac{\Gamma_{\varphi}\tilde{k}^{2}}{2H_{e}\tilde{a}_{\mbox{\tiny RH}}^{1/2}\tilde{m}_{\varphi}^{2}}\right)$ $({\tilde{k}}>{\tilde{m}}_{\varphi}{\tilde{a}}_{\mbox{\tiny RH}})\ .$ (19)
The derivation is discussed in further detail in Appendix D. Lastly, we note that in Region I, there is a discrepancy between the numerical scaling of ${\tilde{k}}^{-6}$ and the exponentially decaying behavior obtained from applying the Stokes phenomenon.
Late reheating ($\tilde{a}_{\mbox{\tiny RH}}>\tilde{m}^{-2/3}$); Conformal coupling $(\xi=1/6)$
$\tilde{m}<1;\ {\tilde{k}}_{\star}={\tilde{m}}^{1/3}$
Region Numerical Inflaton Stokes Steepest Boundary
Scattering Descent Matching
O $2\rightarrow 2$: $\tilde{k}^{-3/2}$ $c{\tilde{m}}^{4}\tilde{k}^{-3/2}$
$1<{\tilde{k}}$
I $\tilde{k}^{3}e^{-\sqrt{\tilde{k}^{3}/\tilde{m}}}$ $\dfrac{\tilde{k}^{3}}{2\pi^{2}}e^{-c_{s}\sqrt{\tilde{k}^{3}/\tilde{m}}}$ $\dfrac{\tilde{k}^{3}}{18}e^{-c_{d}\sqrt{\tilde{k}^{3}/\tilde{m}}}$
${\tilde{k}}_{\star}<\tilde{k}<1$
II $\tilde{k}^{2}$ $c_{m}{\tilde{m}}^{1/3}\tilde{k}^{2}$
$\tilde{m}<\tilde{k}<{\tilde{k}}_{\star}$
III
$0<\tilde{k}<\tilde{m}$
$\tilde{m}>1$
O $2\rightarrow 2$: $\tilde{k}^{-3/2}$ $c{\tilde{m}}^{4}\tilde{k}^{-3/2}$
$2<{\tilde{k}}$ $4\rightarrow 2$: $\tilde{k}^{-15/2}$ $\tilde{k}^{-15/2}$
IV Transition
$\tilde{m}<\tilde{k}<2$
V $\tilde{k}^{2.7}e^{-2.2\tilde{m}}$ $\dfrac{\tilde{k}^{3}}{18}e^{-2\sqrt{3}\tilde{m}}$
[…]